### v1

# Optimal error analysis of an interior penalty virtual element method for fourth-order singular perturbation problems

Abstract
In recent studies [13, 9], the Interior Penalty Virtual Element Method (IPVEM) has been developed for solving a fourth-order singular perturbation problem, with uniform convergence established in the lowest-order case concerning the perturbation parameter. However, the resulting uniform convergence rate is only of half-order, which is suboptimal. In this work, we demonstrate that the proposed IPVEM in fact achieves optimal and uniform error estimates, even in the presence of boundary layers. The theoretical results are substantiated through extensive numerical experiments, which confirm the validity of the error estimates and highlight the method’s effectiveness for singularly perturbed problems.
keywords:
Interior penalty virtual element method; Fourth-order singular perturbation problem; Optimal convergence rate
00footnotetext: The authors are listed in alphabetical order. All authors contributed equally to this paper.
1 Introduction
Let $\Omega$ be a bounded polygonal domain of $\mathbb{R}^{2}$ with boundary $\partial\Omega$. For $f\in L^{2}(\Omega)$, we consider the following boundary value problem of the fourth-order singular perturbation equation:
$\begin{cases}\varepsilon^{2}\Delta^{2}u-\Delta u=f&{\rm in}~~\Omega,\\ u=\dfrac{\partial u}{\partial\boldsymbol{n}}=0&{\rm on}~~\partial\Omega,\end{cases}$ (1.1)
where $\bm{n}=(n_{1},n_{2})$ is the unit outer normal to $\partial\Omega$, and $\varepsilon$ is a real parameter satisfying $0<\varepsilon\leq 1$. In this article, we are primarily concerned with the case where $\varepsilon\to 0$, or the differential equation formally degenerates to the Poisson equation, but with the “extra” normal boundary condition, which will produce boundary layers in the first derivatives of $u$.
This paper focuses on Virtual Element Methods (VEMs), a generalization of the standard finite element method that accommodates general polytopal meshes [2, 1, 3]. Recently, the interior penalty technique for VEMs has been explored in [14] for the biharmonic equation, using the same degrees of freedom (DoFs) as for the $H^{1}$-conforming virtual elements [2]. This numerical scheme can be viewed as a combination of the virtual element space and a discontinuous Galerkin method, as the resulting global discrete space is not $C^{0}$-continuous and an interior penalty formulation is adopted to enforce the continuity of the solution. The interior penalty method is then applied to tackle the fourth-order singular perturbation problem in [13], with certain adaptations to the original space. These modifications include changes in the definition of the $H^{1}$-type projection and the selection of the DoFs. In contrast to the approach in [13], we in [9] adopt the original formulation introduced in [14] to address the fourth-order singular perturbation problem (1.1). To streamline the implementation, we modify the jumps and averages in the penalty term by incorporating the elliptic projector $\Pi_{h}^{\nabla}$.
Both analyses in [13, 9] demonstrate that the IPVEM is robust with respect to the perturbation parameter in the lowest-order case. However, the uniform convergence rate is only $\mathcal{O}(h^{1/2})$, which is suboptimal. To achieve optimal convergence, we can apply the widely used Nitsche’s technique, a method commonly employed for parameter-dependent boundary value problems to enhance performance. Nitsche’s technique weakly enforces boundary conditions, imposing fewer restrictions on the boundary of the domain [8]. Upon reviewing this approach, we find that the IPVEM can be viewed as an application of Nitsche’s technique, which should yield the optimal convergence rate. Building on this insight, we show in this article that the IPVEM in [9] naturally achieves optimal convergence for fourth-order singularly perturbed problems with boundary layers.
2 The continuous variational problem
We first introduce some notations. For a bounded Lipschitz domain $D$ of dimension d(d=1,2)d~(d=1,2), the symbol $(v,w)_{D}=\int_{D}vw\mathrm{d}x$ denotes the $L^{2}$-inner product on $D$, $\|\cdot\|_{0,D}$ or $\|\cdot\|_{D}$ denotes the $L^{2}$-norm. If $D=\Omega$, we abbreviate $\|\cdot\|_{\Omega}$ as $\|\cdot\|$. For a vectorial functions $\bm{v}=(v_{1},v_{2})^{T}$ and $\bm{w}=(w_{1},w_{2})^{T}$, the inner product is defined as $(\bm{v},\bm{w})_{D}=\int_{D}(v_{1}w_{1}+v_{2}w_{2})\mathrm{d}x$. The set of scaled monomials $\mathbb{M}_{r}(D)$ is given by $\mathbb{M}_{r}(D):=\{(\frac{\boldsymbol{x}-\boldsymbol{x}_{D}}{h_{D}})^{\boldsymbol{s}},~~|\boldsymbol{s}|\leq r\}$, with the generic monomial denoted by $m_{D}$, where $h_{D}$ is the diameter of $D$, $\bm{x}_{D}$ is the centroid of $D$, and $r$ is a non-negative integer. For the multi-index ${\boldsymbol{s}}\in{\mathbb{N}^{d}}$, we follow the usual notation $\boldsymbol{x}^{\boldsymbol{s}}=x_{1}^{s_{1}}\cdots x_{d}^{s_{d}}$ with $|\boldsymbol{s}|=s_{1}+\cdots+s_{d}$. Conventionally, $\mathbb{M}_{r}(D)=\{0\}$ for $r\leq-1$. Let $\mathcal{T}_{h}$ denote a tessellation of $\Omega$ into polygons. For a generic element $K$, define $h=\max_{K\in\mathcal{T}_{h}}h_{K}$ and $h_{K}={\rm diam}(K)$. The set $\mathcal{E}_{h}$ represents all the edges in $\mathcal{T}_{h}$. Let $e\subset\partial K$ be the common edge for elements $K=K_{-}$ and $K_{+}$, and let $v$ be a scalar function defined on $e$. We introduce the jump and average of $v$ on $e$ by $v]=v^{-}-v^{+}$ and $\{v\}=\frac{1}{2}(v^{-}+v^{+})$, where $v^{-}$ and $v^{+}$ are the traces of $v$ on $e$ from the interior and exterior of $K$, respectively. On a boundary edge, $[v]=v$ and $\{v\}=v$. For a non-negative integer $m$ and an element $K\in\mathcal{T}_{h}$, denote by $\mathbb{P}_{m}(K)$ the set of all polynomials on $K$ with the total degree no more than $m$. The broken Sobolev space is denoted as
$H^{m}(\mathcal{T}_{h}):=\{w\in L^{2}(\Omega):w|_{K}\in H^{m}(K),K\in\mathcal{T}_{h}\}$
with the broken $H^{m}$-norm
$\|w\|_{m,h}:=(\sum_{K\in\mathcal{T}_{h}}\|w\|_{m,K}^{2})^{1/2},$
and the broken $H^{m}$-seminorm
$|w|_{m,h}=(\sum_{K\in\mathcal{T}_{h}}|v|_{m,K}^{2})^{1/2}.$
Moreover, for any two quantities $a$ and $b$, $a\lesssim b$” indicates $a\leq Cb$” with the constant $C$ independent of the mesh size $h_{K}$, and $a\eqsim b$” abbreviates $a\lesssim b\lesssim a$”. The variational formulation of (1.1) is to find $u\in V:=H_{0}^{2}(\Omega)$ such that
$\varepsilon^{2}a(u,v)+b(u,v)=(f,v),\quad v\in H_{0}^{2}(\Omega),$ (2.1)
where $a(u,v)=(\nabla^{2}u,\nabla^{2}v)$ and $b(u,v)=(\nabla u,\nabla v)$.
For the geometric subdivisions, we make the following assumption (cf. [6, 7]):
  1. 1.
    For each $K\in{\mathcal{T}}_{h}$, there exists a “virtual triangulation” ${\mathcal{T}}_{K}$ of $K$ such that ${\mathcal{T}}_{K}$ is uniformly shape regular and quasi-uniform. The corresponding mesh size of ${\mathcal{T}}_{K}$ is proportional to $h_{K}$. Each edge of $K$ is a side of a certain triangle in ${\mathcal{T}}_{K}$.
Based on this usual geometric assumptions, by the standard Dupont-Scott theory [5], for all $v\in H^{l}(K)$, there exists a certain $q\in\mathbb{P}_{l-1}(K)$ such that
$|v-q|_{m,K}\lesssim h_{K}^{l-m}|v|_{l,K},\quad 0\leq m\leq l.$ (2.2)
3 The interior penalty virtual element method
The interior penalty virtual element method (IPVEM) was first proposed in [14]. In the construction, the authors first introduce a $C^{1}$-continuous virtual element space ($k\geq 2$)
$\widetilde{V}_{k+2}(K)=\left\{v\in H^{2}(K):\Delta^{2}v\in\mathbb{P}_{k}(K),v|_{e}\in\mathbb{P}_{k+2}(e),\partial_{\bm{n}}v|_{e}\in\mathbb{P}_{k+1}(e),~~e\subset\partial K\right\}.$
For the equipped degree of freedoms (DoFs), we refer the reader to [14, 9]. Given $v_{h}\in\widetilde{V}_{k+2}(K)$, Ref. [14] proposes an approximate $H^{1}$-elliptic projection $\Pi_{K}^{\nabla}v_{h}\in\mathbb{P}_{k}(K)$ described by the following equations:
$\left\{\begin{aligned} (\nabla\Pi_{K}^{\nabla}v_{h},\nabla q)_{K}&=-(v_{h},\Delta q)_{K}+\sum\limits_{e\subset\partial K}Q_{2k-1}^{e}(v_{h}\partial_{\bm{n}_{e}}q),\quad q\in\mathbb{P}_{k}(K),\\ \sum_{z\in\mathcal{V}_{K}}\Pi_{K}^{\nabla}v_{h}(z)&=\sum_{z\in\mathcal{V}_{K}}v_{h}(z),\end{aligned}\right$.
where $\mathcal{V}_{K}$ is the set of the vertices of $K$, $Q_{2k-1}^{e}v:=|e|\sum\limits_{i=0}^{k}\omega_{i}v(\bm{x}_{i}^{e})\approx\int_{e}v(s)\mathrm{d}s$, with $(\omega_{i},\bm{x}_{i}^{e})$ being the $(k+1)$ Gauss-Lobatto quadrature weights and points, $\bm{x}_{0}^{e}$ and $\bm{x}_{k}^{e}$ are the endpoints of $e$. In this case, $\Pi_{K}^{\nabla}v_{h}$ is computable by only using the DoFs of the $H^{1}$-conforming virtual element space given by
$V_{k}^{1,c}(K):=\left\{v\in H^{1}(K):\Delta v\in\mathbb{P}_{k-2}(K),~~v|_{\partial K}\in\mathbb{B}_{k}(\partial K)\right\},$
and
$\mathbb{B}_{k}(\partial K)=\{v\in C^{0}(\partial K):v|_{e}\in\mathbb{P}_{k}(e),~~e\subset\partial K\}.$
By using the standard enhancement technique [1], which substitutes the redundant DoFs of $v$ with those of $\Pi_{K}^{\nabla}v$, the local interior penalty virtual element space in [14] is defined as
$V_{k}(K)$ $=\{v\in\widetilde{V}_{k+2}(K):(v,m_{K})_{K}=(\Pi_{K}^{\nabla}v,m_{K})_{K},~m_{K}\in\mathbb{M}_{k}(K)\backslash\mathbb{M}_{k-2}(K),\nabla v(z)=\nabla\Pi_{K}^{\nabla}(z),$
$~\mbox{$z\in\mathcal{V}_{K}$ is any vertex of $K$},(m_{e},\partial_{\bm{n}_{e}}v)_{e}=(m_{e},\partial_{\bm{n}_{e}}\Pi_{K}^{\nabla}v)_{e},~m_{e}\in\mathbb{M}_{k-1}(e),\,e\subset\partial K\}.$
The associated DoFs are given by
  • 1.
    the values of $v(z)$, $z\in\mathcal{V}_{K}$,
  • 2.
    the values of $v(\boldsymbol{x}_{i}^{e})$, $i=1,\cdots,k-1$, $e\subset\partial K$,
  • 3.
    the moments $|K|^{-1}(m_{K},v)_{K}$, $m_{K}\in\mathbb{M}_{k-2}(K)$.
Furthermore, we use $V_{h}$ to denote the global nonconforming virtual element space, which is defined elementwise and required that the DoFs are continuous through the interior vertices and edges while zero for the boundary DoFs.
As usual, we can define the $H^{2}$-projection operator $\Pi_{K}^{\Delta}:V_{k}(K)\to\mathbb{P}_{k}(K)$ by finding the solution $\Pi_{K}^{\Delta}v\in\mathbb{P}_{k}(K)$ of
$\begin{cases}a^{K}(\Pi_{K}^{\Delta}v,q)=a^{K}(v,q),\quad q\in\mathbb{P}_{k}(K),\\ \widehat{\Pi_{K}^{\Delta}v}=\widehat{v},\quad\widehat{\nabla\Pi_{K}^{\Delta}v}=\widehat{\nabla v}\end{cases}$
for any given $v\in V_{k}(K)$, where the quasi-average $\widehat{v}$ is defined by
$\widehat{v}=\frac{1}{|\partial K|}\int_{\partial K}v\mathrm{d}s.$
Lemma 3.1.
[14, 9] Let $m=1,2$ and $0\leq\ell\leq m$. For all $v_{h}\in V_{k}(K)$, there hold
$|v_{h}|_{m,K}\lesssim h_{K}^{\ell-m}|v_{h}|_{\ell,K},$ (3.1)
$|\Pi_{K}^{\nabla}v_{h}|_{m,K}\lesssim|v_{h}|_{m,K},$ (3.2)
$|v_{h}-\Pi_{K}^{\nabla}v_{h}|_{\ell,K}\lesssim h_{K}^{\ell-m}|v_{h}|_{m,K}.$ (3.3)
For the construction of the IPVEM, we do not use the form presented in [14], but instead adopt the modified version from [9], which simplifies the implementation process. For the fourth-order term, the discrete bilinear form is defined by
$a_{h}^{K}(v,w)=a^{K}(\Pi_{K}^{\Delta}v,\Pi_{K}^{\Delta}w)+S^{K}(v-\Pi_{K}^{\Delta}v,w-\Pi_{K}^{\Delta}w),\quad v,w\in V_{k}(K),$
with
$S^{K}(v-\Pi_{K}^{\Delta}v,w-\Pi_{K}^{\Delta}w)=h_{K}^{-2}\sum\limits_{i=1}^{n_{K}}\chi_{i}(v-\Pi_{K}^{\Delta}v)\chi_{i}(w-\Pi_{K}^{\Delta}w),$
where $\{\chi_{i}\}$ are the local DoFs on $K$ with $n_{K}$ being the number of the DoFs. For the second-order term, we define the local bilinear form as
$b_{h}^{K}(v,w):=(\nabla\Pi_{K}^{\nabla}v,\nabla\Pi_{K}^{\nabla}w)_{K}+\sum_{i=1}^{n_{K}}\chi_{i}(v-\Pi_{K}^{\nabla}v)\chi_{i}(w-\Pi_{K}^{\nabla}w).$
The interior penalty virtual element methods for the problem (1.1) can be described as follows: Find $u_{h}\in V_{h}$ such that
$\varepsilon^{2}\mathcal{A}_{h}(u_{h},v_{h})+b_{h}(u_{h},v_{h})=\langle f_{h},v_{h}\rangle,\quad v_{h}\in V_{h},$ (3.4)
where
$\mathcal{A}_{h}(u_{h},v_{h})=a_{h}(u_{h},v_{h})+J_{1}(u_{h},v_{h})+J_{2}(u_{h},v_{h})+J_{3}(u_{h},v_{h}),$
with
$a_{h}(v,w)=\sum_{K\in\mathcal{T}_{h}}a_{h}^{K}(v,w),\quad\,b_{h}(v,w)=\sum_{K\in\mathcal{T}_{h}}b_{h}^{K}(v,w),$
and
[…]
$J_{2}(v,w)=-\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big\{\frac{\partial^{2}\Pi_{h}^{\nabla}v}{\partial\bm{n}_{e}^{2}}\Big\}\Big[\frac{\partial\Pi_{h}^{\nabla}w}{\partial\bm{n}_{e}}\Big]\mathrm{d}s,$
$J_{3}(v,w)=-\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big\{\frac{\partial^{2}\Pi_{h}^{\nabla}w}{\partial\bm{n}_{e}^{2}}\Big\}\Big[\frac{\partial\Pi_{h}^{\nabla}v}{\partial\bm{n}_{e}}\Big]\mathrm{d}s,$
The right-hand side is $\langle f_{h},v_{h}\rangle=\sum\limits_{K\in\mathcal{T}_{h}}\int_{K}f\,\Pi_{0,K}^{k}v_{h}\mathrm{d}x$, where $\Pi_{0,K}^{k}$ is the $L^{2}$ projector onto $\mathbb{P}_{k}(K)$. In what follows, we define
$\|w\|_{h}^{2}=|w|_{2,h}^{2}+J_{1}(w,w),\quad\|w\|_{\varepsilon,h}^{2}=\varepsilon^{2}\|w\|_{h}^{2}+|w|_{1,h}^{2}.$
4 The optimal and uniform error estimate
We first present some results for the error analysis. Since $V_{h}$ and $V_{h}^{1,c}$ share the same DoFs, as done in [9], we can introduce the interpolation from $V_{h}$ to $V_{h}^{1,c}$. For any given $v_{h}\in V_{h}$, let $I_{h}^{c}v_{h}$ be the nodal interpolant of $v_{h}$ in $V_{h}^{1,c}$. According to Lemmas 4.2 and 4.3 in [9], we have the following abstract Strang-type lemma.
Lemma 4.1.
Assume that $u\in H_{0}^{2}(\Omega)\cap H^{k+1}(\Omega)$ is the solution to (2.1) with $k\geq 2$. Then we have the error estimate
$\|u-u_{h}\|_{\varepsilon,h}\lesssim\|u-I_{h}u\|_{\varepsilon,h}+\|u-u_{\pi}\|_{\varepsilon,h}+\|f-f_{h}\|_{V_{h}^{\prime}}+E_{h}(u),$ (4.1)
where $I_{h}$ is the interpolation operator to $V_{h}$ and $u_{\pi}$ is a piecewise polynomial with degree $\leq k$ that satisfies (2.2),
$\|f-f_{h}\|_{V_{h}^{\prime}}=\sup_{v_{h}\in V_{h}\backslash\{0\}}\frac{\langle f-f_{h},v_{h}\rangle}{\|v_{h}\|_{\varepsilon,h}}$
and the consistency term
$E_{h}(u)=\sup_{v_{h}\in V_{h}\backslash\{0\}}\frac{E_{A}(u,v_{h})+E_{J}(u,v_{h})}{\|v_{h}\|_{\varepsilon,h}},$ (4.2)
where
$E_{A}(u,v_{h})=E_{A1}(u,v_{h})+E_{A2}(u,v_{h})+E_{A3}(u,v_{h}),$
$E_{J}(u,v_{h})=\varepsilon^{2}(J_{1}(I_{h}u,v_{h})+J_{3}(I_{h}u,v_{h})),$
with
$E_{A1}(u,v_{h})=\varepsilon^{2}\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big\{\frac{\partial^{2}(u-\Pi_{h}^{\nabla}I_{h}u)}{\partial\boldsymbol{n}_{e}^{2}}\Big\}\left[\frac{\partial v_{h}}{\partial\boldsymbol{n}_{e}}\right]\mathrm{d}s,$
$E_{A2}(u,v_{h})=\varepsilon^{2}\sum_{e\in\mathcal{E}_{h}}\int_{e}\frac{\partial^{2}u}{\partial\boldsymbol{n}_{e}\partial\boldsymbol{t}_{e}}\left[\frac{\partial v_{h}}{\partial\boldsymbol{t}_{e}}\right]\mathrm{d}s,$
$E_{A3}(u,v_{h})=-\varepsilon^{2}\sum_{K\in\mathcal{T}_{h}}(\nabla\Delta u,\nabla v_{h})_{K}+\sum_{K\in\mathcal{T}_{h}}(\nabla u,\nabla I_{h}^{c}v_{h})_{K}-(f,v_{h}).$
Here, $\bm{t}_{e}$ is a unit tangential vector of an edge $e\in\mathcal{E}_{h}$.
The following trace inequalities are very useful for our forthcoming analysis [5, 4].
Lemma 4.2.
For any $K\in\mathcal{T}_{h}$, there hold
$\|v\|_{0,\partial K}\lesssim\|v\|_{0,K}^{1/2}\|v\|_{1,K}^{1/2},\quad v\in H^{1}(K),$ (4.3)
$\|v\|_{0,\partial K}\lesssim h_{K}^{1/2}|v|_{1,K}+h_{K}^{-1/2}\|v\|_{0,K},\quad v\in H^{1}(K).$ (4.4)
As shown in Lemma 3.11 of [14], we can derive the following interpolation error estimate.
Lemma 4.3.
For any $v\in H^{\ell}(K)$ with $2\leq\ell\leq k+1$, it holds
$|v-I_{h}v|_{m,K}\lesssim h_{K}^{\ell-m}|v|_{\ell,K},\quad m=0,1,2.$ (4.5)
Let $u^{0}$ be the solution of the following boundary value problem:
$\begin{cases}-\Delta u^{0}=f&\quad{\rm in}\quad\Omega,\\ u^{0}=0&\quad{\rm on}\quad\partial\Omega.\end{cases}$ (4.6)
The following regularity is well-known and can be found in [10] for instance.
Lemma 4.4.
If $\Omega$ is a bounded convex polygonal domain and $f\in L^{2}(\Omega)$, then
$|u|_{2}+\varepsilon|u|_{3}\lesssim\varepsilon^{-1/2}\|f\|\qquad\text{and}\qquad|u-u^{0}|_{1}\lesssim\varepsilon^{1/2}\|f\|.$
Now we consider the optimal and uniform error estimate. According to Lemma 4.1, it suffices to estimate the right-hand side of (4.1).
Lemma 4.5.
Suppose that $\Omega$ is a bounded convex polygonal domain and that $u^{0}\in H^{m}(\Omega)$ with $2\leq m\leq k+1$. Under the condition of Lemma 4.1, there hold
$\|u-I_{h}u\|_{\varepsilon,h}+\|u-u_{\pi}\|_{\varepsilon,h}\lesssim\varepsilon^{1/2}\|f\|+h^{m-1}|u^{0}|_{m},\quad\|f-f_{h}\|_{V_{h}^{\prime}}\lesssim h^{k}|f|_{k-1}.$
Proof.
For the first term $\|u-I_{h}u\|_{\varepsilon,h}$, we write it explicitly as
$\|u-I_{h}u\|_{\varepsilon,h}^{2}=\varepsilon^{2}|u-I_{h}u|_{2,h}^{2}+\varepsilon^{2}J_{1}(u-I_{h}u,u-I_{h}u)+|u-I_{h}u|_{1,h}^{2}.$
By Lemmas 4.3 and 4.4, we derive
$\varepsilon^{2}|u-I_{h}u|_{2,h}^{2}\leq\varepsilon^{2}\sum_{K\in\mathcal{T}_{h}}|u-I_{K}u|_{2,K}^{2}\lesssim\varepsilon^{2}|u|_{2}^{2}\lesssim\varepsilon\|f\|^{2}.$
Then, apply the trace inequality, the boundedness of $\Pi_{K}^{\nabla}$ in Lemma 3.1, the interpolation error estimate in (4.5) and the regularity estimate in Lemma 4.4 to get
$\varepsilon^{2}J_{1}(u-I_{h}u,u-I_{h}u)$ $=\varepsilon^{2}\sum_{e\in\mathcal{E}_{h}}\int_{e}\frac{\lambda_{e}}{|e|}\Big[\frac{\partial\Pi_{h}^{\nabla}(u-I_{h}u)}{\partial\bm{n}_{e}}\Big]\Big[\frac{\partial\Pi_{h}^{\nabla}(u-I_{h}u)}{\partial\bm{n}_{e}}\Big]\mathrm{d}s$
$\leq\varepsilon^{2}\sum_{k\in\mathcal{T}_{h}}\Big(|\Pi_{h}^{\nabla}(u-I_{h}u)|_{2,K}^{2}+h_{K}^{-2}|\Pi_{h}^{\nabla}(u-I_{h}u)|_{1,K}^{2}\Big)$
$\lesssim\varepsilon^{2}|u|_{2}^{2}\lesssim\varepsilon\|f\|^{2}.$
From Lemma 4.4 and (4.5), we have
$|u-I_{h}u|_{1,h}^{2}$ $=\sum_{K\in\mathcal{T}_{h}}|u-I_{K}u|_{1,K}^{2}=\sum_{K\in\mathcal{T}_{h}}|u-u_{0}+u_{0}-I_{K}u_{0}+I_{K}u_{0}-I_{K}u|_{1,K}^{2}$
$\lesssim\sum_{K\in\mathcal{T}_{h}}\Big(|u-u^{0}-I_{K}(u-u^{0})|_{1,K}^{2}+|u^{0}-I_{K}u^{0}|_{1,K}^{2}\Big)$
$\lesssim\sum_{K\in\mathcal{T}_{h}}\Big(|u-u^{0}|_{1,K}^{2}+h_{K}^{2m-2}|u^{0}|_{m,K}^{2}\Big)$
$\lesssim|u-u_{0}|_{1}^{2}+h^{2m-2}|u^{0}|_{m}^{2}$
$\lesssim\varepsilon\|f\|^{2}+h^{2m-2}|u^{0}|_{m}^{2}.$
Combining above inequalities, we derive
$\|u-I_{h}u\|_{\varepsilon,h}\lesssim\varepsilon^{1/2}\|f\|+h^{m-1}|u^{0}|_{m}.$ (4.7)
Similarly, we can bound the second term as $\|u-u_{\pi}\|_{\varepsilon,h}\lesssim\varepsilon^{1/2}\|f\|+h^{m-1}|u^{0}|_{m}.$ For the third term, from
$\langle f-f_{h},v_{h}\rangle$ $=\sum_{K\in\mathcal{T}_{h}}(f,v_{h}-\Pi_{0,K}^{k}v_{h})_{K}$
$=\sum_{K\in\mathcal{T}_{h}}(f-\Pi_{0,K}^{k-2}f,v_{h}-\Pi_{0,K}^{k}v_{h})_{K}$
$\lesssim\sum_{K\in\mathcal{T}_{h}}h_{K}^{k-1}|f|_{k-1,K}h_{K}|v_{h}|_{1,K}$
$\lesssim h^{k}|f|_{k-1}\|v_{h}\|_{\varepsilon,h},$ (4.8)
we get $\|f-f_{h}\|_{V_{h}^{\prime}}\lesssim h^{k}|f|_{k-1}$. The proof is completed. ∎
The remaining part is to estimate the consistency term $E_{h}(u)$. By Lemma 4.1, we need to consider the estimates of $E_{A1}(u,v_{h})$, $E_{A2}(u,v_{h})$, $E_{A3}(u,v_{h})$, $\varepsilon^{2}J_{1}(I_{h}u,v_{h})$ and $\varepsilon^{2}J_{3}(I_{h}u,v_{h})$.
Lemma 4.6.
Suppose that $\Omega$ is a bounded convex polygonal domain. Under the condition of Lemma 4.1, there holds $E_{h}(u)\lesssim\varepsilon^{1/2}\|f\|$.
Proof.
For $E_{A1}$, by the trace inequality, the inverse inequality of polynomials, the boundedness of $\Pi_{h}^{\nabla}$, (3.1), (4.5) and (2.2), we have
$E_{A1}(u,v_{h})$
$=\varepsilon^{2}\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big\{\frac{\partial^{2}(u-\Pi_{h}^{\nabla}I_{h}u)}{\partial\boldsymbol{n}_{e}^{2}}\Big\}\left[\frac{\partial v_{h}}{\partial\boldsymbol{n}_{e}}\right]\mathrm{d}s\lesssim\varepsilon^{2}\sum_{e\in\mathcal{E}_{h}}\Big\|\frac{\partial^{2}(u-\Pi_{h}^{\nabla}I_{h}u)}{\partial\boldsymbol{n}_{e}^{2}}\Big\|_{0,e}\Big\|\frac{\partial v_{h}}{\partial\boldsymbol{n}_{e}}\Big\|_{0,e}$
$\lesssim\varepsilon^{2}\sum_{e\in\mathcal{E}_{h}}\Big\|\frac{\partial^{2}(u-\Pi_{h}^{\nabla}I_{h}u)}{\partial\boldsymbol{n}_{e}^{2}}\Big\|_{0,e}(h_{K}^{-1/2}|v_{h}|_{1,K}+h_{K}^{1/2}|v_{h}|_{2,K})$
$\lesssim\varepsilon^{2}\sum_{K\in\mathcal{T}_{h}}\Big(h_{K}^{1/2}|u-\Pi_{h}^{\nabla}I_{h}u|_{3,K}+h_{K}^{-1/2}|u-\Pi_{h}^{\nabla}I_{h}u|_{2,K}\Big){(h_{K}^{-1/2}|v_{h}|_{1,K})}$
$\lesssim\varepsilon^{2}\sum_{K\in\mathcal{T}_{h}}\Big(|u-u_{\pi}|_{3,K}+|\Pi_{h}^{\nabla}(u_{\pi}-I_{h}u)|_{3,K}+h_{K}^{-1}|u-u_{\pi}|_{2,K}+h_{K}^{-1}|\Pi_{h}^{\nabla}(u_{\pi}-I_{h}u)|_{2,K}\Big)|v_{h}|_{1,K}$
$\lesssim\varepsilon^{2}\sum_{K\in\mathcal{T}_{h}}\Big(|u-u_{\pi}|_{3,K}+h_{K}^{-1}|u_{\pi}-I_{h}u|_{2,K}+h_{K}^{-1}|u-u_{\pi}|_{2,K}+h_{K}^{-1}|u_{\pi}-I_{h}u|_{2,K}\Big)|v_{h}|_{1,K}$
$\lesssim\varepsilon^{2}\sum_{K\in\mathcal{T}_{h}}|u|_{3,K}|v_{h}|_{1,K}\lesssim\varepsilon^{2}|u|_{3}|v_{h}|_{1,h}\lesssim\varepsilon^{1/2}\|f\||v_{h}|_{1,h}.$
For $E_{A2}$, let $w|_{e}:=\frac{\partial^{2}u}{\partial\boldsymbol{n}_{e}\partial\boldsymbol{t}_{e}}$. Using the trace inequality, the inverse inequality, the projection error estimate and Lemma 4.4, we get
$E_{A2}(u,v_{h})$ $=\varepsilon^{2}\sum_{e\in\mathcal{E}_{h}}\int_{e}\frac{\partial^{2}u}{\partial\boldsymbol{n}_{e}\partial\boldsymbol{t}_{e}}\left[\frac{\partial v_{h}}{\partial\boldsymbol{t}_{e}}\right]\mathrm{d}s\lesssim\varepsilon^{2}\sum_{e\in\mathcal{E}_{h}}\int_{e}(w-\Pi_{0,K}^{k-2}w)\left[\frac{\partial v_{h}}{\partial\boldsymbol{t}_{e}}\right]\mathrm{d}s$
$\lesssim\sum_{e\in\mathcal{E}_{h}}\|w-\Pi_{0,K}^{k-2}w\|_{e}\Big\|\left[\frac{\partial v_{h}}{\partial\boldsymbol{t}_{e}}\right]\Big\|_{e}$
$\lesssim\varepsilon^{2}\sum_{K\in\mathcal{T}_{h}}|u|_{3,K}|v_{h}|_{1,K}\lesssim\varepsilon^{1/2}\|f\||v_{h}|_{1,h}.$
For $E_{A3}$, the integration by parts in (2.1) gives
$\varepsilon^{2}(\nabla(\Delta u),\nabla\varphi)+(\nabla u,\nabla\varphi)=(f,\varphi),\quad\varphi\in C_{0}^{\infty}(\Omega),$
which implies
$-\varepsilon^{2}\sum_{K\in\mathcal{T}_{h}}(\nabla(\Delta u),\nabla I_{h}^{c}v_{h})_{K}+\sum_{K\in\mathcal{T}_{h}}(\nabla u,\nabla I_{h}^{c}v_{h})_{K}=(f,I_{h}^{c}v_{h}),$
since $C_{0}^{\infty}(\Omega)$ is dense in $H_{0}^{1}(\Omega)$, $u\in H^{3}(\Omega)$ and $I_{h}^{c}v_{h}\in H_{0}^{1}(\Omega)$. Then we can write $E_{A3}$ as
$E_{A3}(u,v_{h})=\sum_{K\in\mathcal{T}_{h}}\varepsilon^{2}\big(\nabla\Delta u,\nabla(I_{h}^{c}v_{h}-v_{h})\big)_{K}+(f,I_{h}^{c}v_{h}-v_{h}).$
According to Lemma 4.4 and the error estimate for interpolation operator $I_{h}^{c}$ [9], one gets
$E_{A3}(u,v_{h})$ $\lesssim\sum_{K\in\mathcal{T}_{h}}(\varepsilon^{2}|u|_{3,K}|I_{h}^{c}v_{h}-v_{h}|_{1,K}+h_{K}^{k-1}|f|_{k-1,K}\|I_{h}^{c}v_{h}-v_{h}\|_{0,K})$
$\lesssim\varepsilon^{1/2}\|f\||v_{h}|_{1,h}+h^{k}|f|_{k-1}|v_{h}|_{1,h}.$ (4.9)
For $\varepsilon^{2}J_{1}(I_{h}u,v_{h})$, we apply the trace inequality, the boundedness of $\Pi_{h}^{\nabla}$, (4.5) and (2.2) to get
$\varepsilon^{2}J_{1}(I_{h}u,v_{h})$ $=\varepsilon^{2}\sum_{e\in\mathcal{E}_{h}}\frac{\lambda_{e}}{|e|}\int_{e}\Big[\frac{\partial\Pi_{h}^{\nabla}I_{h}u}{\partial\bm{n}_{e}}\Big]\Big[\frac{\partial\Pi_{h}^{\nabla}v_{h}}{\partial\bm{n}_{e}}\Big]\mathrm{d}s=\varepsilon^{2}\sum_{e\in\mathcal{E}_{h}}\frac{\lambda_{e}}{|e|}\int_{e}\Big[\frac{\partial(\Pi_{h}^{\nabla}I_{h}u-u)}{\partial\bm{n}_{e}}\Big]\Big[\frac{\partial\Pi_{h}^{\nabla}v_{h}}{\partial\bm{n}_{e}}\Big]\mathrm{d}s$
$\leq\varepsilon^{2}\Big(\sum_{e\in\mathcal{E}_{h}}\frac{\lambda_{e}}{|e|}\Big\|\Big[\frac{\partial(\Pi_{h}^{\nabla}I_{h}u-u)}{\partial\bm{n}_{e}}\Big]\Big\|_{0,e}^{2}\Big)^{1/2}J_{1}(v_{h},v_{h})^{1/2}$
$\lesssim\varepsilon\Big(\sum_{K\in\mathcal{T}_{h}}|u-I_{h}u|_{2,K}^{2}+|u-u_{\pi}|_{2,K}^{2}+h_{K}^{-2}(|u-I_{h}u|_{1,K}^{2}+|u-u_{\pi}|_{1,K}^{2})\Big)^{1/2}\|v_{h}\|_{\varepsilon,h}$
$\lesssim\varepsilon|u|_{2}|v_{h}|_{\varepsilon,h}\lesssim\varepsilon^{1/2}\|f\|\|v_{h}\|_{\varepsilon,h}.$
For $\varepsilon^{2}J_{3}(I_{h}u,v_{h})$, the continuity of $u$, the trace inequality, the inverse inequality for polynomials, the boundedness of $\Pi_{h}^{\nabla}$ in (3.2) and the Cauchy-Schwarz inequality give
$\varepsilon^{2}J_{3}(I_{h}u,v_{h})$ $=\varepsilon^{2}\sum_{e\in\mathcal{T}_{h}}\int_{e}\Big[\frac{\partial(\Pi_{h}^{\nabla}I_{h}u-u)}{\partial\bm{n}_{e}}\Big]\Big[\frac{\partial^{2}\Pi_{h}^{\nabla}v_{h}}{\partial\bm{n}_{e}^{2}}\Big]\mathrm{d}s$
$\lesssim\varepsilon^{2}\sum_{e\in\mathcal{T}_{h}}\Big\|\Big[\frac{\partial(\Pi_{h}^{\nabla}I_{h}u-u)}{\partial\bm{n}_{e}}\Big]\Big\|_{0,e}\Big\|\Big[\frac{\partial^{2}\Pi_{h}^{\nabla}v_{h}}{\partial\bm{n}_{e}^{2}}\Big]\Big\|_{0,e}$
$\lesssim\varepsilon^{2}\sum_{K\in\mathcal{T}_{h}}\Big(h_{K}^{1/2}|\Pi_{h}^{\nabla}I_{h}u-u|_{2,K}+h_{K}^{-1/2}|\Pi_{h}^{\nabla}I_{h}u-u|_{1,K}\Big)\Big(h_{K}^{-1/2}|\Pi_{h}^{\nabla}v_{h}|_{2,K}+h_{K}^{1/2}|\Pi_{h}^{\nabla}v_{h}|_{3,K}\Big)$
$\lesssim\varepsilon^{2}\sum_{K\in\mathcal{T}_{h}}\Big(|u-I_{h}u|_{2,K}+|u-u_{\pi}|_{2,k}+h_{K}^{-1}(|u-I_{h}u|_{1,k}+|u-u_{\pi}|_{1,K})\Big){|v_{h}|_{2,K}}$
$\lesssim\varepsilon|u|_{2}\|v_{h}\|_{\varepsilon,h}\lesssim\varepsilon^{1/2}\|f\|\|v_{h}\|_{\varepsilon,h}.$
The proof is completed. ∎
In summary of the above estimates, we conclude that
Theorem 4.1.
Let $u\in H_{0}^{2}(\Omega)\cap H^{k+1}(\Omega)$ be the solution to (2.1) with $k\geq 2$. Suppose that $\Omega$ is a bounded convex polygonal domain and that $u^{0}\in H^{m}(\Omega)$ with $2\leq m\leq k+1$. Then, it follows
$\|u-u_{h}\|_{\varepsilon,h}\lesssim\varepsilon^{1/2}\|f\|+h^{m-1}|u^{0}|_{m}+h^{k}|f|_{k-1}.$ (4.10)
Compared to the result in [9], we obtain the second-order convergence rate in the lowest-order case ($k=2$) if $f\in H^{k-1}(\Omega)$, which is optimal, as $\varepsilon$ approaches zero.
5 Numerical examples
In this section, we report the performance with several examples by testing the accuracy and the robustness with respect to the singular parameter $\varepsilon$ to resolve solutions with boundary layer effects. Unless otherwise specified, we only consider the lowest-order element ($k=2$) and the domain $\Omega$ is taken as the unit square $(0,1)^{2}$.
All examples are implemented in MATLAB R2019b. The code is available on GitHub (https://github.com/Terenceyuyue/mVEM) as part of the mVEM package, which includes efficient and easy-to-follow implementations for various VEMs from the literature. The subroutine ‘Fourth_order_Singular_Perturbation_IPVEM.m’ is used to compute the numerical solutions, while the test script ‘main_Fourth_order_Singular_Perturbation_IPVEM.m’ is used to verify the convergence rates.
Example 5.1.
The solution with a layer is constructed such that $u(x,y)=\varepsilon(\exp(-x/\varepsilon)+\exp(-y/\varepsilon))-x^{2}y$. The right-hand side $f(x,y)=2y$ is independent of the singular parameter $\varepsilon$.
Let $u$ be the exact solution of (1.1), and $u_{h}$ the discrete solution obtained from the underlying VEM (3.4). Since the VEM solution is not explicitly available within the polygonal elements, we will assess the errors by comparing the exact solution $u$ with the elliptic projections. In this manner, the discrete error, measured in terms of the discrete energy norm, is quantified by
$\text{Err}=\Big(\sum\limits_{K\in\mathcal{T}_{h}}(\varepsilon^{2}|u-\Pi_{h}^{\Delta}u_{h}|_{2,K}^{2}+|u-\Pi_{h}^{\nabla}u_{h}|_{1,K}^{2})\Big)^{1/2}/\|f\|.$ (5.1)
To evaluate the accuracy of the proposed method, we analyze a series of meshes based on a Centroidal Voronoi Tessellation of the unit square, with polygonal counts of 32, 64, 128, 256, and 512. These meshes are generated using the MATLAB toolbox PolyMesher [12]. The convergence orders of the errors with respect to the mesh size $h$ are presented in Table 1. It is clear that for small $\varepsilon$, a quadratic convergence rate is observed, which is consistent with the theoretical prediction in Theorem 4.1.
Tab. 1: The convergence rate for Example 5.1
$\N\varepsilon\backslash$ N 32 64 128 256 512 Rate
1e-6 6.4009e-03 3.2484e-03 1.3883e-03 7.1140e-04 3.6204e-04 2.10
1e-8 6.4009e-03 3.2484e-03 1.3883e-03 7.1140e-04 3.6204e-04 2.10
Example 5.2.
In this example, the exact solution takes on a more intricate form, defined as
$u=\left(\exp(\sin\pi x)-1-\pi\varepsilon\frac{\cosh\frac{1}{2\varepsilon}-\cosh\frac{2x-1}{2\varepsilon}}{\sinh\frac{1}{2\varepsilon}}\right)\times\left(\exp(\sin\pi y)-1-\pi\varepsilon\frac{\cosh\frac{1}{2\varepsilon}-\cosh\frac{2y-1}{2\varepsilon}}{\sinh\frac{1}{2\varepsilon}}\right).$
We evaluate the performance on an L-shaped domain (refer to Fig. 1 for a snapshot of the numerical and exact solutions) with $k=3$. The meshes are composed of 100, 200, 300, 400, and 500 polygons, with the corresponding results presented in Table 2. We observe a third-order convergence rate for small $\varepsilon$, consistent with the theoretical prediction in Theorem 4.1.
Tab. 2: The convergence rate for Example 5.2 on an L-shaped domain
$\N\varepsilon\backslash$ N 100 200 300 400 500 Rate
1e-6 1.1531e-04 3.8054e-05 2.0368e-05 1.3161e-05 9.7283e-06 3.13
1e-8 1.1531e-04 3.8054e-05 2.0368e-05 1.3161e-05 9.7283e-06 3.13
Fig. 1: Numerical and exact solutions for Example 5.2 ($\varepsilon=10^{-10}$)
Example 5.3.
For the Poisson problem (4.6), the exact solution is taken as $u^{0}(x,y)=\sin(\pi x)\sin(\pi y)$. Then we compute the right-hand side of (1.1) by $f=-\Delta u^{0}=2\pi^{2}\sin(\pi x)\sin(\pi y)$. The solution $u$ of the singular perturbation problem displays strong boundary layers for very small $\varepsilon$, but the solution $u$ is unknown. So we focus on the error [11]: $\|u^{0}-u_{h}\|_{\varepsilon,h}\lesssim(\varepsilon^{1/2}+h^{2})\|f\|$.
In this example, we repeat the numerical simulation using distorted meshes (cf. [9]). For very small $\varepsilon$, we can ensure a second-order convergence for the error $\|u^{0}-u_{h}\|_{\varepsilon,h}$. By assessing the relative error in (5.1) with $u$ replaced by $u^{0}$, we consistently observe second-order convergence in this example (see Table 3).
Tab. 3: The convergence rate for Example 5.3
$\N\varepsilon\backslash$ N 32 64 128 256 512 Rate
1e-8 8.6803e-03 4.3553e-03 2.2699e-03 1.1487e-03 5.7240e-04 1.95
1e-10 8.6803e-03 4.3553e-03 2.2699e-03 1.1487e-03 5.7240e-04 1.95
Acknowledgements
FF was partially supported by the National Natural Science Foundation of China (NSFC) grant 12401528 and the Fundamental Research Funds for the Central Universities, No. 30924010837. YY was supported by NSFC grant No. 12301561, the Key Project of Scientific Research Project of Hunan Provincial Department of Education (No. 24A0100), the Science and Technology Innovation Program of Hunan Province (No. 2025RC3150) and the 111 Project (No. D23017).